A solution is made by adding 12.7 g of calcium sulfate to 502 g of water. What is the percent mass of calcium sulfate in this solution.
Answer #1

Mass of calcium sulphate =12.7g

Mass of water. = 502g

Total mass of calcium sulfate + water. = 12.7 +502 = 514.7g

Percent mass of calcium sulfate. = 12.7×100/514.7= 2.5%